OpenStudy (anonymous):

rename each decimal as a fraction in simplest form 0.15,0.62

OpenStudy (anonymous):

0.15= 15/100=3/20 0.62=62/100=31/50

OpenStudy (anonymous):

got it ??

OpenStudy (anonymous):

thanks